OpenStudy (anonymous):

Find dy/dx in terms of x and y if 7 x y + 4 x + y = 12

OpenStudy (anonymous):

(7x)(dy/dx) + 7y + 4 + dy/dx = 0 (7x)(dy/dx) + dy/dx = -7y-4 Factor out dy/dx, you get dy/dx = (-7y-4)/(7x+1)

OpenStudy (anonymous):

your welcome.